OpenStudy (anonymous):

2. Factor f(x) = x4 + x3 – 8x2 + 6x + 36 completely. Then sketch the graph.

OpenStudy (anonymous):

you have to consider the possible rational roots. they will be the positive and negative factors of the constant. then you want to use synthetic division with the factors until you get a remainder of 0. when you find a 0 remainder you know that is a factor. then the numbers left from division make a cubic polynomial. rinse repeat

OpenStudy (anonymous):

(x+3)(x+2) = (x^2 + 5x + 6)

OpenStudy (anonymous):

are those your two zeros that you have so far?

OpenStudy (anonymous):

wait (x+2)(x+3)(x2-4x+6)

OpenStudy (anonymous):

that what i got

OpenStudy (anonymous):

but i dont know how to graph it

OpenStudy (anonymous):

you can factor your quadratic term as well

OpenStudy (anonymous):

do you know what the quadratic formula is?

OpenStudy (anonymous):

ax2 + bx + c = 0

OpenStudy (anonymous):

no here is the quadratic formula \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

your quadratic factor. put it into this equation for your last zeros

OpenStudy (anonymous):

x^4+x^3-8x^2+6x+36 =x^4 + 2x^3 - x^3 - 2x^2 - 6x^2 -12x + 18x + 36 = x^3(x+2) - x^2(x+2) - 6x(x+2) + 18(x+2) = (x+2) (x^3 - x^2 - 6x + 18) = (x+2) (x^3 + 3x^2 - 4x^2 - 12x + 6x + 18) = (x+2) {x^2(x+3) - 4x(x+3) + 6(x+3)} = (x+2) (x+3) (x^2-4x+6)

OpenStudy (anonymous):

i didnt check any of that. but i will help you with the next step \[\frac{ 4\pm \sqrt{(-4)^2-4(1)(6)} }{ 2(1) }\] solve this and you have your next zeros

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

get as far as you can and ill help you through step by step. you can do the simple arithmatic easily im sure. do you know how to use i yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

first. solve everything under the root. have you gotten that far?

OpenStudy (anonymous):

sorry i am not good in math

OpenStudy (anonymous):

alight. so (-4)^2 is 16 and -4(6) is -24. so 16-24=-8 correct so you have \[\frac{ -4\pm \sqrt{-8} }{ 2 }\] not you have to use imaginary numbers to get the square root of -8

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

do you know how to simplify radicals?

OpenStudy (anonymous):

no but do u time 4 and 8

OpenStudy (anonymous):

no you will have \[\frac{ -4\pm \sqrt{8}i }{ 2 }\] then you simplify the radical \[\frac{ -4\pm2\sqrt{2} }{ 2 }\] then you can factor out a 2 to get rid of the denominator

OpenStudy (anonymous):

dont forget the i like i did though

OpenStudy (anonymous):

u lost me

OpenStudy (anonymous):

hm... if you want to graph it just make a t table and just graph it. wherever the graph crosses the x axis is a zero. i dont think i will be able to explain it much further without further confusing you. unless you want me to give it a shot

OpenStudy (anonymous):